Grade 12Physical Chemistry

The pH of 0.004M hydrazine solution is 9.7. Calculate its ionization constant Kb and pKb ?

sudhanshu
12 Years agoGrade 12

1 Answer

Supravat
8 Years ago
NH2NH2 + H2O -----> NH2NH3+ + OH-From the given pH the Hion concentration can be measured.So, we have[H+] = antilog (–pH) = antilog (–9.7) = 1.67 × 10^-10Now, [OH-] = Kw / [H+] = 1 × 10^-14/1.67 × 10^-10 = 5.98 × 10^-5.The concentration of the corresponding hydrazine ion is also the same as that of hydroxyl ion. The concentration of both these ions is very small so the concentration of the undissociated base can be taken equal to 0.004 MThus, Kb= [NH2NH3+][OH-] / [NH2NH2]= (5.98 × 10^-5)^2/ 0.004 = 8.96 × 10^-7.So, pKb = –logKb = –log(8.96 × 10–7) = 6.04.